# Bellman Optimality of Average-Reward Robust Markov Decision Processes with a Constant Gain

### v2

@@ 1 Introduction
In this paper, we consider optimality conditions for average-reward robust Markov decision processes (MDPs). Specifically, we consider the finite-state and finite-action setting and we focus on
@@
where $\Pi$ and $\mathrm{K}$ denote the controller’s and adversary’s policy classes, respectively, and $\mu$ is the initial distribution. We consider S-rectangular adversary policy classes, in which the adversary’s perturbation of the transition probabilities is state-wise separable: the choice at one state does not affect the set of admissible choices at any other state. However, action dependencies within each state may remain coupled, thereby enforcing constraints among the adversary’s choices for different actions at the same state. Moreover, both controller and adversary’s policy classes may be history-dependent (denoted by $\Pi_{\mathrm{H}}$ and $\mathrm{K}_{\mathrm{H}}$) or stationary (denoted by $\Pi_{\mathrm{S}}$ and $\mathrm{K}_{\mathrm{S}}$), with potentially asymmetric information structures between the controller and adversary decisions.
@@
$u^{*}(s)=\sup_{\phi\in\mathcal{Q}}\inf_{p_{s}\in\mathcal{P}_{s}}\mathbb{E}_{\phi,p_{s}}\!\left[r(s,A_{0})-\alpha^{*}+u^{*}(X_{1})\right],$ (1.1)
where the expectation is taken w.r.t. the measure $P_{\phi,p_{s}}(A_{0}=a,X_{1}=s^{\prime})=\phi(a)p_{s,a}(s^{\prime})$. Here, $\mathcal{Q}\subseteq\mathcal{P}(\mathcal{A})$ denotes the controller’s admissible decision set at state $s$. In particular, we consider two types of decision sets: (i) the deterministic policy set $\mathcal{Q}=\{\delta_{a}:a\in\mathcal{A}\}$, where $\delta_{a}$ is the Dirac measure at action $a$; and (ii) the fully randomized action set $\mathcal{Q}=\mathcal{A}$. Moreover, $\mathcal{P}_{s}$ denotes the adversary’s decision set, which can be understood as the projection of the ambiguity set $\mathcal{P}$ onto state $s$.
@@
    Weak communication-type assumptions are much harder to deploy and analyze under the controller-adversary dynamics. In particular, different stationary controller/adversary policies can induce different communicating classes. Hence, unlike classical MDPs, it is unclear a priori what the communicating class associated with an optimal policy should be for robust MDPs.
@@ 1.1 Literature review
Robust MDPs: While the Bellman optimality of discounted-reward robust MDPs has been extensively studied [10, 28, 29, 24], the corresponding results for the average-reward setting remain underexplored. To the best of our knowledge, Wang et al., 2023c [26] provides the first results under the stringent assumptions of SA-rectangularity and uniform unichains. Grand-Clement et al., [7] focus on Blackwell optimality, showing that $\epsilon$-Blackwell optimal policies always exist under SA-rectangularity. Moreover, they demonstrate that in S-rectangular RMDPs, average-reward optimal policies may fail to exist; and even when they do exist, they may need to be strictly history-dependent.
@@
Stochastic games (SGs): S-rectangular robust MDPs can be viewed as a generalization of two-player zero-sum SGs. They extend the standard SGs framework in the following ways: (i) an asymmetry of information, where the controller may use history-dependent policies while the adversary is restricted to stationary or Markovian ones, and (ii) ambiguity-set constraints, where the adversary’s feasible set may be infinite and nonconvex, in contrast to the convex mixed-strategy sets of SGs. Below, we provide a detailed review of what is known in the stochastic game literature.
@@
Mertens and Neyman, [15] establish that $\epsilon$-optimal strategies always exist for all players, implying that every zero-sum stochastic game admits a value in the finite-state and finite-action setting. In the special case of irreducible stochastic games, Section 5 of Filar and Vrieze, [5] shows that both players possess optimal stationary strategies, that the optimal value is state-independent, and that the Bellman equation admits a solution equal to this common value. A concise overview of these foundational results can also be found in the tutorial by Renault, [19]. To further weaken these assumptions, Wei et al., [27] show that when the SG is player 1 communicating, the Bellman equation still admits a solution that characterizes the value of the game. The player 1 communication condition is equivalent to the controller-communication assumption in our paper. However, as in the literature on stochastic games, their players’ policies can be randomized, making the decision sets convex. For continuous-state or continuous-action settings, analogous results have been derived under the geometric ergodicity assumption [13, 11, 12].
@@
Finally, Garrec, [6] study communicating zero-sum product stochastic games, where each player has an individual state that evolves solely according to their own previous state and action. However, their notion of communication differs from the one considered in our formulation. In fact, their setting does not satisfy the (weakly) communicating assumption used in this paper. As shown by Garrec, [6] as well as Vigeral, [23], Sorin and Vigeral, [21], Ziliotto, [31], the value of average-reward stochastic games may fail to exist without compactness or communicating assumptions when the action spaces are infinite.
@@ 2 Canonical Construction and the Optimal Robust Control Problem
We also define the random element $H_{t}:\Omega\rightarrow\mathbf{H}_{t}$ by point evaluation $H_{t}(\omega)=h_{t}$, and the $\sigma$-field $\mathcal{H}_{t}:=\sigma(H_{t})$.
@@ 3.1 Robust Bellman Equation with a Constant Gain
It is useful to introduce the following discounted robust Bellman equation. These will serve as key theoretical tools in establishing the existence of solutions to the average-reward equation (3.1).
@@
$v_{\gamma}^{*}:S\rightarrow\mathbb{R}$ is the unique solution of the $\gamma$-discounted robust Bellman equation if
@@
Similarly, $v^{\prime}_{\gamma}:S\rightarrow\mathbb{R}$ is the unique solution to the $\gamma$-discounted inf-sup equation if
@@
Remark.
In the discounted setting, existence and uniqueness of solutions to (3.3) and (3.4) follow from a standard contraction mapping argument; see Wang et al., 2023a [24]. By contrast, in the average-reward setting, existing results establish the existence of a solution to (3.1) only in SA-rectangular settings and under additional assumptions [26, Theorem 8]. We substantially generalize these existence results to one–sided weak communication settings, which are robust analogues of the weakly communicating structures commonly assumed in classical MDPs to ensure constant-gain optimality [18].
@@ 3.3 Optimality of Stationary Policies From the Robust Bellman Equation
Similar to the classical MDP setting, stationary policy extracted from the robust Bellman equation are $\epsilon$-optimal for any $\epsilon>0$.
@@
Let $(u^{*},\alpha^{*})$ be a solution to (3.1). If for some $\epsilon\geq 0$ and $\Delta:S\rightarrow\mathcal{Q}$,
@@ 4.2 Weakly Communicating Structures
In this section, we establish that under compactness and the one-sided weakly communicating structures in Definition 3, both the robust Bellman equation (3.1) and the inf–sup equation (3.2) admit solutions. These one-sided weak-communication assumptions are well motivated: they are extensively used in the classical MDP literature to guarantee a constant optimal gain [18]. In particular, as reviewed earlier, one-sided communication ensures the existence of solutions in SG settings where both players may employ fully randomized strategies [27], which in turn induces convex controller policy and adversary ambiguity sets. However, to the best of our knowledge, extensions to the weak-communication setting have not been established in the literature.
@@
  • A stationary controller policy $\Delta:S\rightarrow\mathcal{Q}$ is communicating if it is weakly communicating with $C_{\Delta}=S$.
    The controller is communicating if every stationary policy $\Delta:S\rightarrow\mathcal{Q}$ is communicating.
  • A stationary adversary policy $p\in\mathcal{P}$ is communicating if it is weakly communicating with $C_{p}=S$.
    The adversary is communicating if every stationary policy $p\in\mathcal{P}$ is communicating.
Remark.
Note that our (weak) communication definitions parallel their classical MDP counterpart. Moreover, a controller/adversary may be (weakly) communicating even when the communicating sets $C_{\Delta}$$C_{p}$ depend on the particular stationary policy; they need not coincide across stationary policies.
@@
Symmetric to the controller-side results, we show that the adversary-side weak communication structures will imply the existence of solutions to (3.2). The proof for Theorem 6 is deferred to Appendix D.
@@
Note that Theorem 6 does not, by itself, guarantee a solution to (3.1). Nevertheless, under suitable sufficient conditions, the solution $(u^{\prime},\alpha^{\prime})$ of (3.2) also satisfies (3.1). Let $\delta_{a}\in\mathcal{P}(\mathcal{A})$ denote the Dirac measure at $a\in A$.
@@ 4.3 Useful Corollaries
Assume that $\mathcal{Q}$ is compact and that for each $\Delta:S\rightarrow\mathcal{Q}$ there exists $p\in\mathcal{P}$ such that $p_{\Delta}$ is irreducible. Then the constant gain average reward robust Bellman equation (3.1) has a solution.
@@
Given this flexibility, Corollary 7.1 could be easy to verify in applications where the ambiguity set is a distributional ball around a nominal kernel $p_{0}\in\mathcal{P}$. In particular, if either (i) $p_{0}$ induces an irreducible Markov chain for every policy, or (ii) $\mathcal{P}$ contains transition probabilities with full support (e.g., total variation, $L^{p}$ [4], and Wasserstein ambiguity sets), then (3.1) admits a solution.
Proof of Corollary 7.1.
By the assumptions in the first claim, for any $\Delta:S\rightarrow\mathcal{Q}$, there is $p\in\mathcal{P}$ so that $p_{\Delta}$ is irreducible. In particular, all states communicate under $p_{\Delta}$; i.e. $\forall s,s^{\prime}\in S$, $p_{\Delta}^{N}(s^{\prime}|s)>0$ for some $N\geq 1$. This verifies the assumption (2) of Theorem 5 and implies the first statement of Corollary 7.1.
For the second statement, the same argument shows that the assumption (2) of Theorem 6 is satisfied. ∎
@@
Proof of Corollary 7.2.
Fix $\Delta:S\rightarrow\mathcal{Q}$. For every $p\in\mathcal{P}$, let $R_{\Delta}(p)\subseteq S$ denote the closed recurrent class of $p_{\Delta}$. We define
@@
Therefore, $\Delta$ is weakly communicating. As $\Delta:S\rightarrow\mathcal{Q}$ is arbitrary, we conclude that the controller is weakly communicating. Thus, the assumption (1) of Theorem 5 is satisfied, implying Corollary 7.2. ∎
@@
Building on Theorems 5 and 6, we establish a more general result, Corollary 7.3, which allows stationary policies to induce chains with multiple recurrent classes.
Before we state the result, for fixed $\Delta:S\rightarrow\mathcal{Q}$ and $p\in\mathcal{P}$, we define the collections of recurrent classes
@@
The proof of Corollary 7.3 is deferred to Appendix E.
@@ 5 Proof of Theorem 5
Our proof primarily addresses the weakly communicating case in assumption (1). Since a communicating controller is a special case, the argument under assumption (2) carries over with only minor changes. Importantly, assumption (2) does not require compactness of $\mathcal{P}_{s}$ for $s\in S$. We will explain how the proof extends under this weaker assumption and why compactness is unnecessary in that case; see Remark 3 and 5.
@@
where the first inequality follows from the construction of $G_{\Delta_{k}}$ in (5.1) and the second inequality is because of (5.3). Also note that by Lemma 1, $N=N^{w,y,\Delta_{k}}\leq|S|$. This proves Lemma 3. ∎
@@
Notice that in the proof of Lemma 3, only the property of $O_{\Delta}$ was used, not that of $K_{\Delta}$. Thus, to establish Lemma 3, it suffices to construct a finite subcover from the collection $O_{\Delta}:\Delta:S\to\mathcal{Q}$}. Since the construction of $O_{\Delta}$ does not require the compactness of $\mathcal{P}_{s}$ for $s\in S$, Lemma 3 holds even without assuming compactness of the adversary’s ambiguity set. The same observation applies to Lemma 4.
@@
So, under assumption (2) of the theorem, we only need to uniformly bound $\xi_{1}$, which does require the compactness of $\mathcal{P}_{s}$ for $s\in S$; see Remark 5.
@@
Note that by the construction of $\kappa$ in (5.6)
@@
Here, $(i)$ is becasue $x\neq y$, $(ii)$ follows from the definition of $E_{x}^{\Delta_{\epsilon},\kappa}$, $(iii)$ is because by (5.6)
@@
On the other hand,
$E^{\Delta_{\epsilon},\kappa}_{x}\sum_{k=\tau_{y}}^{\infty}\gamma^{k-\tau_{y}}r(X_{k},A_{k})$ (5.9)
@@
$\stackrel{{\scriptstyle(iii)}}{{=}}E^{\Delta_{\epsilon},p_{\epsilon}}_{x}\sum_{k=\tau_{y}}^{\infty}\gamma^{k-\tau_{y}}r(X_{k},A_{k})$
@@
$\stackrel{{\scriptstyle(iv)}}{{=}}E^{\Delta_{\epsilon},p_{\epsilon}}_{y}\sum_{k=0}^{\infty}\gamma^{k}r(X_{k},A_{k})$
@@
Here $(i)$ applies Fubini’s theorem leveraging the positivity of the summand, $(ii)$ follows from the definition of $\kappa$ in (5.6), $(iii)$ is obtained from reversing the previous equalities, and $(iv)$ is because $p_{\epsilon}$ is a stationary policy and, under $E_{x}^{\Delta_{\epsilon},p_{\epsilon}}$, $(X_{t},A_{t})$ is strong Markov.

### v3

@@ 1 Introduction
In this paper, we consider optimality conditions for average-reward robust Markov decision processes (MDPs). Specifically, we consider the finite-state and finite-action setting and define
@@
where $\Pi$ and $\mathrm{K}$ denote the controller’s and adversary’s policy classes, respectively, $\mu$ is the initial distribution, and $r(\cdot,\cdot)\in[0,1]$ is a bounded reward function. We consider S-rectangular adversary policy classes, in which the adversary’s perturbation of the transition probabilities is state-wise separable: the choice at one state does not affect the set of admissible choices at any other state. However, action dependencies within each state may remain coupled, thereby enforcing constraints among the adversary’s choices for different actions at the same state. Moreover, both controller and adversary’s policy classes may be history-dependent (denoted by $\Pi_{\mathrm{H}}$ and $\mathrm{K}_{\mathrm{H}}$) or stationary (denoted by $\Pi_{\mathrm{S}}$ and $\mathrm{K}_{\mathrm{S}}$), with potentially asymmetric information structures between the controller and adversary decisions.
@@
$u^{*}(s)=\sup_{\phi\in\mathcal{Q}}\inf_{p_{s}\in\mathcal{P}_{s}}E_{\phi,p_{s}}\!\left[r(s,A_{0})-\alpha^{*}+u^{*}(X_{1})\right],$ (1.1)
where the expectation is taken w.r.t. the measure $P_{\phi,p_{s}}(A_{0}=a,X_{1}=s^{\prime})=\phi(a)p_{s,a}(s^{\prime})$. Here, $\mathcal{Q}\subseteq\mathcal{P}(\mathcal{A})$ denotes the controller’s admissible decision set at state $s$. In particular, we consider two types of decision sets: (i) the deterministic policy set $\mathcal{Q}=\{\delta_{a}:a\in\mathcal{A}\}$, where $\delta_{a}$ is the Dirac measure at action $a$; and (ii) the fully randomized action set $\mathcal{Q}=\mathcal{P}(\mathcal{A})$. Moreover, $\mathcal{P}_{s}$ denotes the adversary’s decision set, which can be understood as the projection of the ambiguity set $\mathcal{P}$ onto state $s$.
@@
    Weak communication-type assumptions are much harder to analyze under the controller-adversary dynamics. In particular, different stationary controller/adversary policies can induce different communicating classes. Hence, unlike classical MDPs, it is unclear a priori what the communicating class associated with an optimal policy should be for robust MDPs.
@@ 1.1 Literature review
Robust MDPs: While the Bellman optimality of discounted-reward robust MDPs has been extensively studied [14, 36, 37, 31, 10], the corresponding results for the average-reward setting remain underexplored. To the best of our knowledge, Wang et al., 2023c [34] provides the first results under strong assumptions of SA-rectangularity and uniform unichains. Grand-Clement et al., [9] focus on Blackwell optimality, showing that $\epsilon$-Blackwell optimal policies always exist under SA-rectangularity. Moreover, they demonstrate that in S-rectangular RMDPs, average-reward optimal policies may fail to exist; and even when they do exist, they may need to be strictly history-dependent.
@@
Stochastic games (SGs): S-rectangular robust MDPs can be viewed as a generalization of two-player zero-sum SGs. They extend the standard SGs framework in the following ways: (i) an asymmetry of information, where the controller may use history-dependent policies while the adversary is restricted to stationary or Markovian ones, and (ii) ambiguity-set constraints, where the adversary’s feasible set may be infinite and nonconvex, in contrast to the convex mixed-strategy sets of SGs. Below, we provide a detailed review of what is known in the stochastic game literature. The properties of Blackwell $\varepsilon$-optimal strategies are subsequently studied in Grand-Clément and Vieille, [11].
@@
Tanaka et al., [29] show that stochastic games have a value and that both players have optimal stationary policies when the Bellman equation admits a solution, which establishes a result similar to our Theorem 1. Mertens and Neyman, [19] establishes that $\epsilon$-optimal strategies always exist for all players, implying that every zero-sum stochastic game admits a value in the finite-state and finite-action setting. In the special case of irreducible stochastic games, Section 5 of Filar and Vrieze, [5] shows that both players possess optimal stationary strategies, that the optimal value is state-independent, and that the Bellman equation admits a solution equal to this common value. A concise overview of these foundational results can also be found in the tutorial by Renault, [24]. To further weaken these assumptions, Wei et al., [35] shows that when the SG is player 1 communicating, the Bellman equation still admits a solution that characterizes the value of the game. The player 1 communication condition is the SG analogue of the controller-communication assumption in our paper. However, as in the literature on stochastic games, their players’ policies are randomized, making the decision sets convex. For continuous-state or continuous-action settings, analogous results have been derived under the geometric ergodicity assumption [17, 15, 16].
@@
Finally, Garrec, [6] study communicating zero-sum product stochastic games, where each player has an individual state that evolves solely according to their own previous state and action. However, their notion of communication differs from the one considered in our formulation. In fact, their setting does not satisfy the (weakly) communicating assumption used in this paper. As shown by Garrec, [6] as well as Vigeral, [30], Sorin and Vigeral, [27], Ziliotto, [40], the value of average-reward stochastic games may fail to exist without compactness or communicating assumptions when the action spaces are infinite. Recently, Gaubert et al., [7] study Blackwell optimality in stochastic games.
@@ 2 Canonical Construction and the Optimal Robust Control Problem
We also define the random element $H_{t}:\Omega\rightarrow\mathbf{H}_{t}$ by $H_{t}(\omega)=h_{t}$, and the $\sigma$-field $\mathcal{H}_{t}:=\sigma(H_{t})$.
@@
Without loss of generality, we assume the reward function $r$ is bounded between 0 and 1.
@@ 3.1 Robust Bellman Equation with a Constant Gain
It is useful to introduce the following discounted robust Bellman equation for a discount factor $\gamma\in(0,1)$. These will serve as key theoretical tools in establishing the existence of solutions to the average-reward equation (3.1).
@@
We say that $v_{\gamma}^{*}:S\rightarrow\mathbb{R}$ solves the $\gamma$-discounted robust Bellman equation if
@@
Similarly, $v^{\prime}_{\gamma}:S\rightarrow\mathbb{R}$ solves the $\gamma$-discounted inf-sup equation if
@@
Remark 1.
In the discounted setting, existence and uniqueness of solutions to (3.3) and (3.4) follow from a standard contraction mapping argument; see Wang et al., 2023a [31]. By contrast, in the average-reward setting, existing results establish the existence of a solution to (3.1) only in SA-rectangular settings and under additional assumptions [34, Theorem 8]. We substantially generalize these existence results to one–sided weak communication settings, which are robust analogues of the weakly communicating structures commonly assumed in classical MDPs to ensure constant-gain optimality [23].
@@ 3.3 Optimality of Stationary Policies From the Robust Bellman Equation
Similar to the classical MDP setting, given $(u^{*},\alpha^{*})$, any stationary policy that is $\epsilon$-optimal for the robust Bellman equation is also $\epsilon$-optimal for the long-run average reward of the robust MDP, for any $\epsilon\geq 0$.
@@
Let $(u^{*},\alpha^{*})$ be a solution to (3.1). If for some $\epsilon\geq 0$ and a stationary policy $\Delta:S\rightarrow\mathcal{Q}$,
@@ 4.2 Weakly Communicating Structures
In this section, we establish that under compactness and the one-sided weakly communicating structures in Definition 3, both the robust Bellman equation (3.1) and the inf–sup equation (3.2) admit solutions. These one-sided weak-communication assumptions are well motivated: they are extensively used in the classical MDP literature to guarantee a constant optimal gain [23]. In particular, as reviewed earlier, one-sided communication ensures the existence of solutions in SG settings where both players may employ fully randomized strategies [35], which in turn induces convex controller policy and adversary ambiguity sets. However, to the best of our knowledge, extensions to the non-convex or the weak-communication setting have not been established in the literature.
@@
Remark 4.
Note that our weak communication definitions parallel their classical MDP counterpart. Moreover, a controller/adversary may be weakly communicating even when the communicating sets $C_{\Delta}$$C_{p}$ depend on the particular stationary policy; they need not coincide across stationary policies.
@@
  • A stationary controller policy $\Delta:S\rightarrow\mathcal{Q}$ is communicating if it is weakly communicating with $C_{\Delta}=S$. The controller is communicating if every $\Delta:S\rightarrow\mathcal{Q}$ is communicating.
  • A stationary adversary policy $p\in\mathcal{P}$ is communicating if it is weakly communicating with $C_{p}=S$. The adversary is communicating if every $p\in\mathcal{P}$ is communicating.
@@
Symmetric to the controller-side results, we show that the adversary-side weak communication structures will imply the existence of solutions to (3.2). The proof for Theorem 6 is deferred to Appendix D as it is similar to the proof of Theorem 5.
@@
Note that Theorem 6 does not, by itself, guarantee a solution to (3.1). Nevertheless, under suitable sufficient conditions, the solution $(u^{\prime},\alpha^{\prime})$ of (3.2) also satisfies (3.1). We state these conditions formally in Theorem 7, where we let $\delta_{a}\in\mathcal{P}(\mathcal{A})$ denote the Dirac measure at $a\in\mathcal{A}$. The proof of Theorem 7 is provided in Appendix E.
@@ 4.3 Useful Corollaries
Assume that $\mathcal{Q}$ is compact and that for each controller’s stationary policy $\Delta:S\rightarrow\mathcal{Q}$ there exists $p\in\mathcal{P}$ such that $p_{\Delta}$ is irreducible. Then the constant gain average reward robust Bellman equation (3.1) has a solution.
@@
Given this flexibility, Corollary 7.1 could be easy to verify in applications where the ambiguity set is a distributional ball around a nominal kernel $p_{0}\in\mathcal{P}$. In particular, if either (i) $p_{0}$ induces an irreducible Markov chain for every policy, or (ii) $\mathcal{P}$ contains transition probabilities with full support (e.g., any S- and SA-rectangular total variation [38], $L^{p}$ [4], and Wasserstein [8] ambiguity sets with a radius $\delta>0$), then (3.1) admits a solution.
Proof.
Proof of Corollary 7.1 By the assumptions in the first claim, for any $\Delta:S\rightarrow\mathcal{Q}$, there is $p\in\mathcal{P}$ so that $p_{\Delta}$ is irreducible. In particular, all states communicate under $p_{\Delta}$; i.e. $\forall s,s^{\prime}\in S$, $p_{\Delta}^{N}(s^{\prime}|s)>0$ for some $N\geq 1$. This verifies the assumption (2) of Theorem 5 and implies the first statement of Corollary 7.1.
For the second statement, the same argument shows that (2) in Theorem 6 is satisfied. ∎∎
@@
Proof.
Proof of Corollary 7.2 Fix $\Delta:S\rightarrow\mathcal{Q}$. For every $p\in\mathcal{P}$, let $R_{\Delta}(p)\subseteq S$ denote the closed recurrent class of $p_{\Delta}$. We define
@@
Therefore, $\Delta$ is weakly communicating. As $\Delta:S\rightarrow\mathcal{Q}$ is arbitrary, we conclude that the controller is weakly communicating. Thus, the assumption (1) of Theorem 5 is satisfied, implying Corollary 7.2. ∎∎
@@
Building on Theorems 5 and 6, we establish a more general result, Corollary 7.3, which allows stationary policies to induce chains with multiple recurrent classes. Before we state the result, for fixed $\Delta:S\rightarrow\mathcal{Q}$ and $p\in\mathcal{P}$, we define the collections of recurrent classes
@@
Here, a closed communicating class is a set of states that all communicate with each other and from which no state can reach any state outside the set [22].
@@
In other words, an overlap-connected family of sets $\mathcal{R}_{\Delta}$ means that for any two sets in the family, they can be reached from one to the other by a finite chain of sets in $\mathcal{R}_{\Delta}$ such that each consecutive pair in the chain has a nonempty intersection. The same interpretation also applies to $\mathcal{R}_{p}$.
@@
The proof of Corollary 7.3 is deferred to Appendix F.
@@ 5 Proof of Theorem 5
Our proof primarily addresses the weakly communicating case in assumption (1). Since a communicating controller is a special case, the argument under assumption (2) carries over with only minor changes. Importantly, assumption (2) does not require compactness of $\mathcal{P}_{s}$ for $s\in S$. We will explain how the proof extends under this weaker assumption and why compactness is unnecessary in that case; see Remark 6 and 8.
@@
where the first inequality follows from the construction of $G_{\Delta_{k}}$ in (5.1) and the second inequality is because of (5.3). Also note that by Lemma 1, $N=N^{w,y,\Delta_{k}}\leq|S|$. This proves Lemma 3. ∎∎
@@
Notice that in the proof of Lemma 3, only the property of $O_{\Delta}$ was used, not that of $K_{\Delta}$. Thus, to establish Lemma 3, it suffices to construct a finite subcover from the collection $\left\{{O_{\Delta}:\Delta\in\left\{{S\to\mathcal{Q}}\right\}}\right\}$. Since the construction of $O_{\Delta}$ does not require the compactness of $\mathcal{P}_{s}$ for $s\in S$, Lemma 3 holds even without assuming compactness of the adversary’s ambiguity set. The same observation applies to Lemma 4.
@@
So, under assumption (2) of the theorem, we only need to uniformly bound $\xi_{1}$, which will not require the compactness of $\mathcal{P}_{s}$ for $s\in S$; see Remark 8.
@@
Note that by the construction of $\kappa$ in (5.6), we have
@@
Here, $(i)$ is because $x\neq y$, $(ii)$ follows from the definition of $E_{x}^{\Delta_{\epsilon},\kappa}$, $(iii)$ is because by (5.6)
@@
On the other hand, we observe
$E^{\Delta_{\epsilon},\kappa}_{x}\sum_{k=\tau_{y}}^{\infty}\gamma^{k-\tau_{y}}r(X_{k},A_{k})$
@@
Here $(i)$ applies Fubini’s theorem leveraging the positivity of the summand, $(ii)$ follows from the definition of $\kappa$ in (5.6). From $(ii)$, reversing the previous equalities, we have that
$E^{\Delta_{\epsilon},\kappa}_{x}\sum_{k=\tau_{y}}^{\infty}\gamma^{k-\tau_{y}}r(X_{k},A_{k})$ $=E^{\Delta_{\epsilon},p_{\epsilon}}_{x}\sum_{k=\tau_{y}}^{\infty}\gamma^{k-\tau_{y}}r(X_{k},A_{k})$ (5.9)
@@
$\stackrel{{\scriptstyle(i)}}{{=}}E^{\Delta_{\epsilon},p_{\epsilon}}_{y}\sum_{k=0}^{\infty}\gamma^{k}r(X_{k},A_{k})$
@@
where $(iv)$ is because $p_{\epsilon}$ is a stationary policy and, under $E_{x}^{\Delta_{\epsilon},p_{\epsilon}}$, $\left\{{X_{t},A_{t}:t\geq 0}\right\}$ is strong Markov.